OpenStudy (anonymous):

How does newton's method work ? This is all supposed to be a review, but I'm having trouble remembering how it works... f(x)=x^3 + x + 3 (Initial Guess = 1)

OpenStudy (anonymous):

\[x_{n+1}=x_{n}-\frac{f(x_n)}{f'(x_n)}\]

OpenStudy (anonymous):

first guess is 1 \(f(1)=5\) so \(1\) is kind of a lousy guess, since you want it to be 0

OpenStudy (anonymous):

\[f'(x)=3x^2+1\] so \[f'(1)=4\]

OpenStudy (anonymous):

second guess is therefore \(x_2=1-\frac{5}{4}=-\frac{1}{4}\)

OpenStudy (anonymous):

lather, rinse repeat it gets real messy real fast

OpenStudy (anonymous):

I'm kind of remembering how to do this

OpenStudy (anonymous):

you can also do some algebra if you like before you start \[x-\frac{f(x)}{f'(x)}=x-\frac{x^3+x+3}{3x^2+1}\] \[=\frac{2x^3-3}{3x^2+1}\] and plug in there, but it amounts to the same thing

OpenStudy (anonymous):

Thank you very much for your help :)

OpenStudy (anonymous):

yw